OpenStudy (anonymous):

Find the length of arc AB given the following: Round your answer to two decimal places. Use 3.14 for pi. y = degree of angle r = radius r=5 y=69

OpenStudy (anonymous):



OpenStudy (anonymous):

i need stpes not the answer necessarily

OpenStudy (campbell_st):

the angle subtended by the arc AB is 69/360 the arc length is l = 69/360 x 2 x pi x 5

OpenStudy (campbell_st):

or this the fraction of the circle, times circumference \[C = 2\pi r\] fraction 69/360 arc length \[69/360 \times 2 \times \pi \times 5\]

OpenStudy (anonymous):

163.91

OpenStudy (campbell_st):

I don't have a calculator... but it seems a bit large... as circumference of the whole circle is 31.4 cm.... then you need about 1/6 of that more like 5.5 or something close to that

OpenStudy (anonymous):

crap thx